OpenStudy (anonymous):

A lighthouse is 200ft from a long straight wall on the beach. The light in the lighthouse rotates through on complete rotation once every 4 sec. Find the equation that gives the distance d in terms of time t. Then, to the nearest tenth, find d when t is 1/3 sec. (It has something to do with velocities).

OpenStudy (anonymous):

You want to find distance in terms of the angle first, Then find the angle in terms of t

OpenStudy (anonymous):

A lighthouse is 200ft from a long straight wall on the beach.  This tells you: \[ d\cos\theta = 200 \implies d = 200\cos(\theta) \]

OpenStudy (anonymous):

Whoops, I mean \[ d = \frac{200}{\cos(\theta)} \]

OpenStudy (anonymous):

Now you need to find \(\theta\) in terms ot \(t\).

OpenStudy (anonymous):

"The light in the lighthouse rotates through on complete rotation once every 4 sec." This means when \(\Delta t=4\), then \(\Delta \theta = 2\pi\)

OpenStudy (anonymous):

Would it be \[D=200\tan \frac{ \pi }{ 2}t \] though?  The drawing looks like this

OpenStudy (anonymous):

Ohhh, so not the distance from the lighthouse to the wall, but just the distance from the center of the wall?

OpenStudy (anonymous):

In that case: \[ \tan(\theta) = \frac{d}{200} \implies d = 200\tan(\theta) \]

OpenStudy (anonymous):

Right next to the 'd' on the left is a right angle by the way, sorry. If that helps.

OpenStudy (anonymous):

We know that \(\theta = ct\) and \(\theta(4)=2\pi = c(4)\) which means \(c = \pi/2\) and \(\theta(t) = \pi t/2\)

OpenStudy (anonymous):

Meaning: \(d = 200\tan(\pi t/2)\)

OpenStudy (anonymous):

Would t be on the outside?

OpenStudy (anonymous):

No, it's in the function. Why would it be on the outside? \(\theta\) was on in inside.